How many gallons of a 10% acid solution and a 30% acid solution need to be mixed to make 20 gallons of a 22% acid solution?
Found 2 solutions by richwmiller, fcabanski:
Answer by richwmiller(17219)

x+y=20
.1x+.3y=.22*20
x=8 y=12

Answer by fcabanski(1391)

Set up two equations: the first for the % solution and the second for the amount (gallons). x = gallons of the first solution, and y = gallons of the second solution.


Gallons: x+y=20 s0 x = 20-y
Solution: .1x +.3y = .22*20 = 4.4 | Substitute the x value from the first eq: 20-y
.1(20-y) + .3y = 4.4
2-.1y+.3y = 4.4
.2y = 2.4
y =12


From the first equation x = 20-y = 20-12 = 8.


8 gallons of the 10% solution and 12 gallons of the 30% solution.


Contact me for private tutoring or private problem solving.